# WCCNet: Wavelet-context Cooperative Network for Efficient Multispectral Pedestrian Detection

### v2

@@ III-B Multi-domain Crossmodal Rearranging Fusion
Multimodal fusion of different modalities is important for accurate multispectral pedestrian detection. In our framework, RGB features $\mathbf{I}_{R}^{j}$ in spatial domain and infrared features $\mathbf{I}_{T}^{j}$ in frequency domain have their own characteristic semantics in channel dimension. The simple amplification for numerically different features in existing approaches[12, 19, 20, 16] may ignore the potential semantic gap between those heterogeneous features of different spectra, and turns out to be nearly ineffective as demonstrated in Sec. IV-D.
In spite of semantic diversity, there exists inherent spatial consistency cross multimodal feature maps. In light of this, we propose a novel fusion module named Crossmodal Rearranging Fusion (CMRF). As illustrated in Fig. 3, CMRF consists of three parts: Cross-domain Embedding (CE) layer, Crossmodal Spatial Alignment (CSA) layer, and Semantic Rearranging Fusion (SRF) layer.
Cross-domain Embedding
@@
At the $j^{th}$ stage, infrared features $\mathbf{I}_{T}^{j}$ contains frequency-domain components which is complementary but contrasting to RGB features $\mathbf{I}_{R}^{j}$. Therefore, Cross-domain Embedding Layer (CE) are adopted to embed $\mathbf{I}_{T}^{j}$ into the same domain of features in RGB branch. The embedding layer consists of a few embedding units, specifically two units in our implementation, which conduct $3\times 3$ convolution, batch normalization, and activation function in sequence. Taking one embedding unit as an example, for the input tensor $\mathbf{I}_{T}^{j}$, the embedding unit generates the output tensor $\mathbf{F}_{T}^{j}\in\mathbb{R}^{C_{T}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ of the same size,
@@
where $\alpha$ and $\beta$ denote learnable affine parameters, and function $\mathcal{F}_{conv}(\cdot)$ denotes the convolution operation with learnable weights. $\mathrm{E}(\cdot)$ and $\sigma(\cdot)$ denote the mean and the standard deviation respectively. Leaky ReLU[40] (LReLU) is adopted as activation. In this way, for further crossmodal fusion, the $j^{th}$-stage CE layer projects $\mathbf{I}_{T}^{j}$ into semantically embedded infrared features $\mathbf{F}_{T}^{j}$.
Crossmodal Spatial Alignment
Since large amounts of multispectral image pairs are weakly aligned as represented in Fig. 4, it is vital to align feature maps of two spectra to assure crossmodal spatial consistency for subsequent crossmodal fusion. Inspired by [41], we propose Crossmodal Spatial Alignment layer to predict pixel-wise offsets of infrared modality relative to RGB modality for alignment. Differing from self-shifting in [41], the offsets in our framework are predicted from features of different modalities simultaneously and utilized for crossmodal shifting.
@@
(a) KAIST Training Set
(b) KAIST Test Set
Semantic Rearranging Fusion
@@
Despite the features of infrared modality are accommodated to RGB modality so far, we argue that the aim of multimodal fusion for multispectral pedestrian detection is to adaptively amplify complementary features of pedestrians which are semantically similar across different modalities. Instead of numerically complementary feature extraction in prior works, we propose Semantic Rearranging Fusion (SRF) layer which leverages inherent crossmodal spatial consistency and transfers semantically complementary infrared features in relevant local regions to complement the corresponding RGB features.
@@
We first briefly explain the overall process. Recall that generic RGB features $\mathbf{I}_{R}^{j}\in\mathbb{R}^{C_{R}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ contain global structures and instance textures, and aligned infrared features $\mathbf{F}_{AT}^{j}\in\mathbb{R}^{C_{T}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ are derived from CSA layer as the input multimodal tensors for SRF. Following crossmodal spatial consistency, SRF exploits the relationship between channel-dimension semantic contents of the pixel $\mathbf{I}_{R}^{j}(l)$ located at $l=(m,n)$ of $\mathbf{I}_{R}^{j}$ and the corresponding local region centered at $l=(m,n)$ in $\mathbf{F}_{AT}^{j}$. The channel-dimension semantic contents of $\mathbf{I}_{R}^{j}(l)$ are utilized to generate the corresponding local rearranging weight $\mathbf{W}_{l}$. Then $\mathbf{W}_{l}$ is applied to extract instance-specific complementary features from $\mathbf{F}_{AT}^{j}$ through calculating the weighted sum for pixels of $\mathbf{F}_{AT}^{j}$ spatially relevant to $\mathbf{I}_{R}^{j}$, namely the pixels of corresponding local region centered at $l$ in $\mathbf{F}_{AT}^{j}$.
@@
To be more specific, convolution layers are first performed on $\mathbf{I}_{R}^{j}$ to compress the channel number $C_{R}^{j}$ into a smaller one $C_{RK}^{j}$, resulting in $\mathbf{I}_{RK}^{j}\in\mathbb{R}^{C_{RK}^{j}\times\frac{H}{2^{j}}\times\frac{H}{2^{j}}}$. Then given $C_{RK}^{j}=k_{r}^{2}$ where $k_{r}\in\mathbb{N}^{*}$, the pixel $\mathbf{I}_{RK}^{j}(l)\in\mathbb{R}^{k_{r}^{2}\times 1\times 1}$ located at $l=(m,n)$ is spatially reshaped to get the local rearranging weight $\mathbf{W}_{l}\in\mathbb{R}^{k_{r}\times k_{r}}$. The obtained $\mathbf{W}_{l}$ is spatially normalized by softmax function to avoid altering the mean of the corresponding region in $\mathbf{F}_{AT}^{j}$, which sets the sum of elements in $\mathbf{W}_{l}$ to $1$. The generation of $\mathbf{W}_{l}$ mentioned above can be formally written as
$\mathbf{W}_{l}=\mathcal{S}\left(\mathcal{F}_{conv}(\mathbf{I}_{R}^{j}(l))\right),$ (10)
@@
where $\mathcal{F}_{conv}$ represents convolution operations with learnable weights, and $\mathcal{S}(\cdot)$ refers to softmax function. It is worth noting that $\mathbf{W}_{l}$ generated from $\mathbf{I}_{R}^{j}(l)$ will give high-response values if $\mathbf{I}_{R}^{j}(l)$ suspected to contain pedestrian instances, which is illustrated by the visualization of $\mathbf{W}_{l}$ in Sec. IV-D.
@@
At last, instance-specific complementary features $\mathbf{G}_{T}^{j}$ of infrared modality are extracted by convolving the corresponding local region in $\mathbf{F}_{AT}^{j}$ with $\mathbf{W}_{l}$ centered at $(m,n)$, which can be expressed as
@@
where $k=\lfloor\frac{k_{r}}{2}\rfloor$. In this way, detailed features $\mathbf{G}_{T}^{j}$ from infrared modality of the same instance with RGB modality are extracted, and then concatenated with $\mathbf{I}_{R}^{j}$ to semantically complement the corresponding RGB features for further pedestrian detection with higher precision.
IV Experiments
In this section, the performance of the proposed WCCNet is evaluated on several important multispectral pedestrian detection benchmarks and compared with other state-of-the-art methods. Comprehensive ablation experiments are also conducted with quantitative and qualitative illustrations.
IV-A Datasets
@@
Our approach is evaluated on two widely-used multispectral pedestrian detection benchmarks including KAIST[5] and FLIR[6], whose data samples are all collected in autonomous driving scenes. Next, we introduce their detailed information.
KAIST
@@
KAIST[5] contains videos of 95,328 RGB-infrared paired frames with 103,128 bounding boxes in total, which are taken from urban road scenes during day and night to cover changes in diverse lighting conditions, heavy occlusions, and multiple pedestrian scales. The videos are split into two halves following [5], where the first half is used for training, and the latter half for testing. Then those two set of videos are sampled every 20 frames into images to obtain image datasets for training and testing respectively. Among images for training, $20$ percent of them are randomly selected for validation in training phase. Following common annotation settings for a fair comparison with recent related works, the paired annotations in [9] are adopted for training and the cleaned annotations[17] are applied for testing. Despite the application of sanitized annotations, there still exist significant spatial misalignments in KAIST as illustrated in Fig. 4, especially for training set, which calls for methods to counter crossmodal misalignments.
FLIR
@@
FLIR[6] contains targets of richer categories than KAIST, including person, car and bicycle. RGB and infrared image pairs in FLIR are taken from moving vehicles at various situations. Following common annotation settings[42], misaligned RGB-infrared image pairs are removed, resulting in well-aligned 4,129 image pairs for training and 1,013 image pairs for testing.
IV-B Implementation Details
TABLE I: Comparisons with the state-of-the-art methods on KAIST dataset. - means not reported in the corresponding paper. >> means the estimated lower bound for the metrics not reported in the corresponding paper. Red, blue, and green colors denote the first, the second, and the third place respectively.
@@ IV-B Implementation Details
Fusion RPN+BF[18] >138.4>138.4 >126.20>126.20 800 - - - - 29.83 30.51 27.62
MSDS[8] >71.56>71.56 >480.03>480.03 228 TITAN X 34.20 32.12 38.83 11.34 10.54 12.94
ARCNN[9] >43.88>43.88 >353.79>353.79 120 1080TI 34.95 34.36 36.12 9.34 9.94 8.38
CIAN[10] >162.5>162.5 >142.19>142.19 66.67 1080TI 35.57 36.06 32.38 14.12 14.78 11.13
TC-DET[11] >63.00>63.00 >156.13>156.13 33.17 1080 - - - 27.11 34.81 10.31
MBNet[12] >99.23>99.23 >139.96>139.96 70 1080TI 31.87 32.39 30.95 8.13 8.28 7.86
MLPD[13] >162.5>162.5 >142.19>142.19 12 2080TI 28.49 28.39 28.69 7.58 7.96 6.95
RISNet[20] - - - - - - - 7.89 7.61 7.08
BAANet[14] >73.96>73.96 - 70 1080TI - - - 7.92 8.37 6.98
CMPD[15] >99.23>99.23 >139.96>139.96 110 1080TI 28.98 28.30 30.56 8.16 8.77 7.31
TINet[16] 61.00 - 87.7 TITAN X - - - 9.15 10.25 7.48
WCCNet 17.50 89.98 10.44 3090 19.05 21.17 13.36 7.37 8.55 4.68
@@ 33.61 TITAN X
WCCNet-L 48.35 231.81 20.84 3090 18.41 21.06 12.76 7.28 8.36 5.08
WCCNet-S 9.75 53.29 6.39 3090 19.37 21.97 13.34 7.56 8.63 4.98
WCCNet-XS 7.78 43.56 4.29 3090 22.85 25.51 16.19 10.78 12.36 7.32
(a) All-dataset setting for the whole day
(b) All-dataset setting in the daytime
(c) All-dataset setting at night
(d) Reasonable setting for the whole day
(e) Reasonable setting in the daytime
(f) Reasonable setting at night
Figure 5: Log-scale MR-FPPI curves on KAIST with different experimental settings (all-dataset setting and reasonable setting) and different illumination conditions. The lower the curve, the higher the accuracy represented.
Platform and Parameters
The proposed WCCNet has extended the framework of YOLOX[32] in PyTorch. Note that ADWT layers excel at extracting complementary infrared features, and can reciprocally work with other multiscale CNN backbones via CMRF fusion modules. We utilize NVIDIA RTX 3090 for training, and employ a series of GPUs for testing, as shown in Table I. WCCNet is trained from scratch by Stochastic Gradient Descent (SGD) for 80 epochs with an initial learning rate of 0.001, a weight decay of 0.02, and a batch size of 12. Other hyperparameters follow YOLOX if not mentioned. For preprocessing, RGB images and infrared images are all resized to $640\times 640$ before being input to network. Besides, Mosaic[43] and MixUp[44] are adopted as data augmentation methods for the first three quarters of training epochs. For post-processing, the threshold of Non-Maximum Suppression (NMS) is $0.65$, and the threshold for confidence score of positive targets is set to $0.01$.
Supplementary Details for Network
A lightweight convolutional residual backbone named DarkNet-34 is proposed for WCCNet, which follows DarkNet[45] family and has 34 convolutional layers. Our DarkNet-34 has five stages with $\{1,2,4,3,2\}$ residual blocks respectively. The first three stages are applied in RGB branch for feature extraction and the last two stages are shared by two modalities for further multispectral detection. The number of channels $C_{R}^{j}$ at $j^{th}$ stage is adjustable with a scaling parameter $\tau$ and set to $2^{j}\times 32\tau$, while the number of channels for $j^{th}$-stage embedded feature maps $F_{T}^{j}$ in CE layer of infrared branch is set to $2^{j}\times 16\tau$. In our implementation, the $\tau$ of WCCNet is set to $0.5$. Besides, for comprehensive analysis on the impacts of different amounts of parameters, we also present three additional types of WCCNet, namely WCCNet-L with $\tau=1.0$, WCCNet-S with $\tau=0.25$, and WCCNet-XS with $\tau=0.125$. Smaller $\tau$ means lower computational cost but greater underfitting risk, resulting in faster inference speed but lower accuracy. WCCNet with $\tau=0.5$ is recommended in practical applications as a trade-off between detection accuracy and speed.
For the neck of WCCNet, feature pyramid mechanisms are adopted, including Spatial Pyramid Pooling (SPP)[46] and Feature Pyramid Network (FPN)[31]. Specifically, SPP is applied after the last stage of the backbone for further extraction of high-level semantic information. Besides, FPN is applied for top-down information interaction between features of different scales. The rest part of WCCNet, namely detection head for classification and regression of bounding boxes, remains consistent with YOLOX.
Quantitative Metrics
The metric $\mathrm{MR}^{-2}$ is used for evaluating the accuracy on KAIST. It equals to the area of log-average Miss Rate (MR) over False Positive Per Image (FPPI), where FPPI ranges from $10^{-2}$ to $1$. Besides, the log-average MR-FPPI curves are obtained by varying the threshold of detector from 0 to $1$ to generate different MR-FPPI points and plotting these points in logarithmic-scale axes. For consistent comparison with previous works on FLIR, the COCO-style[47] evaluation metrics are adopted, namely Average Precision (AP) for specific class and mean Average Precision (mAP) for all classes. To evaluate the efficiency of multispectral pedestrian detectors, the memory cost are denoted by the numbers of parameters (Param.), while computational complexity are indicated by the number of floating point operations (FLOPs) theoretically and the inference time (Time) practically. Note that the inference time of our model includes post-processing time. Many previous works may not prioritize the efficiency of their models and have not report corresponding metrics in their papers, thus we estimate their lower bound. The FLOPs metric is estimated based on $640\times 640$ input size.
IV-C Comparison with State-of-the-arts
Comparison on KAIST
We evaluate WCCNet and conduct comparisons with other state-of-the-art methods [18, 8, 9, 10, 11, 12, 13, 14, 15, 20, 16] under reasonable and all-dataset settings following [5], as illustrated in Table I and Fig 5. The reasonable setting means that the evaluation is conducted on a subset of testing set which only consists of pedestrians taller than 55 pixels without heavy occlusion, while all-dataset setting evaluates detectors on the whole testing set.
Accuracy Comparison. The experimental results on detection precision are given in terms of $\mathrm{MR}^{-2}$. As can be observed in Table I, all types of our proposed WCCNet outperforms previous methods under all-dataset and reasonable settings. Besides, the MR-FPPI curves under those two settings are shown in Fig. 5(b) and Fig. 5(c) respectively, which also demonstrates the superiority of WCCNet. Especially for all-dataset setting which is more challenging, the WCCNet-L has taken a lead of $35.38\%$ percent over the best recorded method MLPD[13] with 18.41 $\mathrm{MR}^{-2}$ on the all-day subset. This demonstrates the strong capability of WCCNet to handle small and occlusion pedestrians. Moreover, for detection in low-illumination conditions, the precision results on the night subset under reasonable setting show that WCCNet surpasses previous methods by a large margin no less than $32.66\%$. Therefore, it confirms that our ADWT-integrated infrared branch can extract sufficient representations of pedestrians in low-illumination conditions with remarkably fewer parameters.
Efficiency Comparison. For evaluation on model efficiency, the number of parameters, FLOPs, and the inference time are provided in Table. I. Besides, the Speed vs. $\mathrm{MR}^{-2}$ curves under reasonable setting are shown in Fig. 5(a). It is obvious that, compared with previous works, WCCNet achieves remarkable fast speed with fewer parameters and lower computational cost while maintaining superior accuracy, which also supports the computational complexity analysis in Sec. III-A. Especially, WCCNet-S with only 9.75MM parameters and 53.29GG FLOPs can infer at a speed of $156.49$ frames per second (FPS) while maintain competitive accuracy, which again shows the efficiency of WCCNet and its potential in practical applications.
Comparison on FLIR
@@ IV-C Comparison with State-of-the-arts
TABLE II: Evaluation results on FLIR dataset. Red, blue, and green colors denote the first, the second, and the third place respectively. - means not reported by the paper of corresponding methods.

### v3

@@ 3.2 Multi-domain Crossmodal Rearranging Fusion
Multimodal fusion of different modalities is important for accurate multispectral pedestrian detection. In our framework, RGB features $\mathbf{I}_{R}^{j}$ in spatial domain and infrared features $\mathbf{I}_{T}^{j}$ in frequency domain have their own characteristic semantics in channel dimension. The simple amplification for numerically different features in existing approaches [MBNet2020, tang2022piafusion, wang2022RISNet, zhang2023TripleNet] may ignore the potential semantic gap between those heterogeneous features of different spectra, and turns out to be nearly ineffective as demonstrated in Sec. 4.5. In spite of semantic diversity, there exists inherent spatial consistency cross multimodal feature maps. In light of this, we propose a novel fusion module named Crossmodal Rearranging Fusion (CMRF). As illustrated in Fig. 4, CMRF consists of three parts: Cross-domain Embedding (CE) layer, Crossmodal Spatial Alignment (CSA) layer, and Semantic Rearranging Fusion (SRF) layer.
3.2.1 Cross-domain Embedding
@@
At the $j^{th}$ stage, infrared features $\mathbf{I}_{T}^{j}$ contains frequency-domain components which is complementary but contrasting to RGB features $\mathbf{I}_{R}^{j}$. Therefore, Cross-domain Embedding Layer (CE) are adopted to embed $\mathbf{I}_{T}^{j}$ into the same domain of features in RGB branch. The embedding layer consists of a few embedding units, specifically two units in our implementation, which conduct $3\times 3$ convolution, batch normalization, and activation function in sequence. Taking one embedding unit as example, for input tensor $\mathbf{I}_{T}^{j}$, the embedding unit generates the output tensor $\mathbf{F}_{T}^{j}\in\mathbb{R}^{C_{T}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ of the same size,
@@
where $\alpha$ and $\beta$ denote learnable affine parameters, and function $\mathcal{F}_{conv}(\cdot)$ denotes the convolution operation with learnable weights. $\mathrm{E}(\cdot)$ and $\sigma(\cdot)$ denote the mean and the standard deviation respectively. Leaky ReLU (LReLU) is adopted as activation. In this way, for further crossmodal fusion, the $j^{th}$-stage CE layer projects $\mathbf{I}_{T}^{j}$ into semantically embedded infrared features $\mathbf{F}_{T}^{j}$.
(a) KAIST Training Set
(b) KAIST Sanitized Test Set
3.2.2 Crossmodal Spatial Alignment
Since multispectral image pairs are often weakly aligned, it is vital to align feature maps of two spectra to assure crossmodal spatial consistency for subsequent crossmodal fusion. Inspired by [dai2017deformable], we propose Crossmodal Spatial Alignment layer to predict pixel-wise offsets of infrared modality relative to RGB modality for alignment. Differing from self-shifting in [dai2017deformable], the offsets in our framework are predicted from features of different modalities simultaneously and utilized for crossmodal shifting.
@@ 3.2.2 Crossmodal Spatial Alignment
3.2.3 Semantic Rearranging Fusion
@@
Despite the features of infrared modality are accommodated to RGB modality so far, the features further need to adaptively amplify complementary information of pedestrians which are semantically similar across different modalities. Instead of numerically complementary feature extraction in prior works, we propose Semantic Rearranging Fusion (SRF) layer for semantic-aware fusion. SRF leverages inherent crossmodal spatial consistency and transfers semantically complementary infrared features in relevant local regions to complement the input multimodal representation, further recalling reliable pedestrian features from infrared modality.
@@
We first briefly explain the overall process. Recall that multimodal representation $\mathbf{F}_{M}^{j}\in\mathbb{R}^{C_{M}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ contain combined pedestrian patterns, and aligned infrared features $\mathbf{F}_{AT}^{j}\in\mathbb{R}^{C_{T}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ are derived from CSA layer as the input multimodal tensors for SRF. Following crossmodal spatial consistency, SRF exploits the relationship between channel-dimension semantic contents of the pixel $\mathbf{F}_{M}^{j}(l)$ located at $l=(m,n)$ of $\mathbf{F}_{M}^{j}$ and the corresponding local region centered at $l=(m,n)$ in $\mathbf{F}_{AT}^{j}$. The channel-dimension semantic contents of $\mathbf{F}_{M}^{j}(l)$ are utilized to generate the corresponding local rearranging weight $\mathbf{W}_{l}$. Then $\mathbf{W}_{l}$ is applied to extract instance-specific complementary features from $\mathbf{F}_{AT}^{j}$ through calculating the weighted sum for pixels of $\mathbf{F}_{AT}^{j}$ spatially relevant to $\mathbf{F}_{M}^{j}$, namely the pixels of corresponding local region centered at $l$ in $\mathbf{F}_{AT}^{j}$.
@@
To be more specific, convolution layers are first performed on $\mathbf{F}_{M}^{j}$ to compress the channel number $C_{M}^{j}$ into a smaller one $C_{MK}^{j}$, resulting in $\mathbf{F}_{MK}^{j}\in\mathbb{R}^{C_{MK}^{j}\times\frac{H}{2^{j}}\times\frac{H}{2^{j}}}$. Then given $C_{MK}^{j}=k_{r}^{2}$ where $k_{r}\in\mathbb{N}^{*}$, the pixel $\mathbf{F}_{MK}^{j}(l)\in\mathbb{R}^{k_{r}^{2}\times 1\times 1}$ located at $l=(m,n)$ is spatially reshaped to get the local rearranging weight $\mathbf{W}_{l}\in\mathbb{R}^{k_{r}\times k_{r}}$. The obtained $\mathbf{W}_{l}$ is spatially normalized by softmax function to avoid altering the mean of the corresponding region in $\mathbf{F}_{AT}^{j}$, which sets the sum of elements in $\mathbf{W}_{l}$ to $1$. The generation of $\mathbf{W}_{l}$ mentioned above can be formally written as
$\mathbf{W}_{l}=\textrm{Softmax}\left(\mathcal{F}_{conv}(\mathbf{F}_{M}^{j}(l))\right),$ (10)
@@
where $\mathcal{F}_{conv}$ represents convolution operations with learnable weights. $\mathbf{W}_{l}$ generated from $\mathbf{I}_{R}^{j}(l)$ will focus on $\mathbf{F}_{M}^{j}(l)$ if $\mathbf{F}_{M}^{j}(l)$ suspected to contain pedestrian instances.
@@
At last, instance-specific complementary features $\mathbf{G}_{T}^{j}$ of infrared modality are extracted by convolving the corresponding local region in $\mathbf{F}_{AT}^{j}$ with $\mathbf{W}_{l}$ centered at $(m,n)$,
@@
where $k=\lfloor\frac{k_{r}}{2}\rfloor$. At last, $\mathbf{G}_{T}^{j}$ is concatenated with $\mathbf{F}_{M}^{j}$ as next stage input, semantically complementing pedestrian features, which is visualized in Sec. 4.5.
4 Experiments
4.1 Benchmarks
@@
Our approach is evaluated on two widely-used multispectral pedestrian detection benchmarks, i.e., KAIST and FLIR.
@@
KAIST [KAIST2015] contains videos of 95,328 RGB-infrared paired frames with 103,128 bounding boxes in total, which are taken from urban road scenes during day and night to cover changes in diverse lighting conditions, heavy occlusions, and multiple pedestrian scales. The videos are split into two halves following [KAIST2015], where the first half is used for training, and the latter half for testing. Then those two set of videos are sampled every 20 frames into images to obtain image datasets for training and testing respectively. Among images for training, $20$ percent of them are randomly selected for validation in training phase. Following common annotation settings for a fair comparison with recent related works, the paired annotations in [ARCNN2019] are adopted for training set and the sanitized annotations [liu2016CleanKAIST] are applied for testing set.
@@
FLIR [FLIRDataset] contains targets of richer categories than KAIST, including person, car and bicycle. RGB and infrared image pairs in FLIR are taken from moving vehicles at various situations. Following common annotation settings [CFR2020], misaligned RGB-infrared image pairs are removed, resulting in 4,129 image pairs for training and 1,013 image pairs for testing.
4.2 Implementation Details
4.2.1 Platform and Data Processing
The proposed WCCNet has extended the framework of YOLOX [ge2021yolox] in PyTorch. We utilize NVIDIA RTX 3090 GPUs for training, and employ a series of GPUs for testing, as shown in Table 1. For preprocessing, RGB images and infrared images are all resized to $640\times 640$ before being input to network. For post-processing, the threshold of Non-Maximum Suppression (NMS) is $0.65$, and the threshold for confidence score of positive targets is set to $0.01$. We split training process into two stages, namely pretraining and supervised training.
Figure 6: Pretraining strategy of WCCNet backbone. Elaborated data augmentation operations are applied.
4.2.2 Pretraining Setting
The backbone of WCCNet is first pretrained on a union dataset to be distilled from DINOv3-ConvNeXt-Base [dinov3] model following [beyer2022knowledge], as shown in Fig. 6. The union dataset combines COCO-Minitrain-25k dataset [COCO2014] and 50% images with Top-50% average target size from KAIST training set. WCCNet backbone is trained by AdamW optimizer for 50 epochs with initial learning rate of 1e-3 and total batch size of 64. Extracted multiscale features from CMRF modules in WCCNet backbone is constrained to be consistent with the ones of DINOv3 by MSE loss, and EMA strategy is adopted.
Images first undertakes random transformation, including random cropping, rotation, color jittering, and etc. Then augmented images are sent to WCCNet and DINOv3 respectively. Notably, for infrared branch of WCCNet, images are converted to grayscale and have 50% probability to be randomly shifted to simulate misalignment. The shift range is set to $\pm 0.1W$ pixels, where $W$ denotes the width of feature maps, and the blank area caused by shifting is padded with zeros.
@@ 4.2.2 Pretraining Setting
Fusion RPN+BF[RPN2017Konig] >138.4>138.4 >126.20>126.20 800 - - - - 29.83 30.51 27.62
MSDS[MSDS2018] >71.56>71.56 >480.03>480.03 228 TITAN X 34.20 32.12 38.83 11.34 10.54 12.94
ARCNN[ARCNN2019] >43.88>43.88 >353.79>353.79 120 1080TI 34.95 34.36 36.12 9.34 9.94 8.38
CIAN[CIAN2019] >162.5>162.5 >142.19>142.19 66.67 1080TI 35.57 36.06 32.38 14.12 14.78 11.13
TC-DET[TCDET2020] >63.00>63.00 >156.13>156.13 33.17 1080 - - - 27.11 34.81 10.31
MBNet[MBNet2020] >99.23>99.23 >139.96>139.96 70 1080TI 31.87 32.39 30.95 8.13 8.28 7.86
MLPD[MLPD2021] >162.5>162.5 >142.19>142.19 12 2080TI 28.49 28.39 28.69 7.58 7.96 6.95
RISNet[wang2022RISNet] - - - - - - - 7.89 7.61 7.08
BAANet[yang2022BAANet] >73.96>73.96 - 70 1080TI - - - 7.92 8.37 6.98
CMPD[CMPD2022] >99.23>99.23 >139.96>139.96 110 1080TI 28.98 28.30 30.56 8.16 8.77 7.31
GAFF[GAFF2021] 31.43 - 9.34 1080TI 27.30 30.59 19.22 6.48 8.35 3.46
TINet[zhang2023TripleNet] 61.00 - 87.7 TITAN X - - - 9.15 10.25 7.48
TFNet[TFNet2024] - - 150 TITAN X - - - 8.09 8.41 7.34
ICAFusion[ICAFusion] 120.2 - 26 3090 - - - 7.17 6.82 7.85
MS-DETR[MS-DETR2024] 76.38 258.84 110 3090 20.64 22.76 15.77 6.13 7.78 3.18
WCCNet 17.50 91.00 10.62 3090 17.81 20.36 12.37 6.72 7.80 4.72
@@ 36.43 TITAN X
WCCNet-S 9.75 53.86 6.56 3090 19.54 21.77 14.39 7.30 8.53 5.03
WCCNet-XS 7.78 43.87 4.45 3090 20.57 22.61 15.68 8.92 10.06 6.62
@@
Table 1: Comparisons with the state-of-the-art methods on KAIST dataset. - means not reported in the corresponding paper. >> means the estimated lower bound for the metrics not reported in the corresponding paper. Red, blue, and green colors denote the first, the second, and the third place respectively.
4.2.3 Supervised Training Setting
WCCNet is further trained on training sets of involved benchmarks. Pretrained backbone weight is loaded for initialization. AdamW optimizer is adopted for 80 epochs with initial learning rate of 1e-3 and total batch size of 32. Mosaic and MixUp data augmentations are adopted for the first three quarters of training epochs.
All-dataset setting
(a) All Time
(b) Daytime
(c) Night
Reasonable setting
(d) All Time
(e) Daytime
(f) Night
Figure 7: Log-scale MR-FPPI curves on KAIST with different experimental settings and different illumination conditions. The lower the curve, the higher the accuracy represented.
4.2.4 Supplementary Details for Network
A lightweight convolutional network is adopted as RGB-branch backbone, which follows DarkNet [redmon2018yolov3] family and has 34 convolutional layers. It has five stages with $\{1,2,4,3,2\}$ residual blocks respectively. The first three stages are applied in RGB branch for feature extraction and the last two stages are shared by two modalities for further multispectral detection. The number of channels $C_{R}^{j}$ at $j^{th}$ stage is adjustable with a scaling parameter $\tau$ and set to $2^{j}\times 16\tau$, while the number of channels for $j^{th}$-stage embedded feature maps $F_{T}^{j}$ in CE layer of infrared branch is set to $2^{j}\times 8\tau$. In our implementation, the $\tau$ of WCCNet is set to $1.0$. Besides, for comprehensive analysis on the impacts of different amounts of parameters, we also present two more efficient types of WCCNet, namely WCCNet-S with $\tau=0.5$, and WCCNet-XS with $\tau=0.25$. Smaller $\tau$ means lower computational cost but greater underfitting risk, resulting in faster inference speed but lower accuracy.
For MoWE, routed expert number $N_{e}$ is set to 8, activated expert number $N_{act}$ is set to 4, and shared expert number $N_{s}$ is set to 2. Haar and db3 wavelet are selected for static DWT experts, and the support length $K$ of ADWT experts is set to 6. Pooling size for MoWE router is set to $4\times 4$.
For the neck of WCCNet, feature pyramid mechanisms are adopted, including Spatial Pyramid Pooling (SPP) [SPP2015He] and Feature Pyramid Network (FPN) [lin2017FPN]. Specifically, SPP is applied after the last stage of the backbone for further extraction of high-level semantic information. Besides, FPN is applied for top-down information interaction between features of different scales. Detection head for classification and regression of bounding boxes are following YOLOX [ge2021yolox].
Figure 8: Comparison on detection results with prior works[MBNet2020, MLPD2021, GAFF2021, MS-DETR2024], which are overlaid on RGB/infrared image pairs in two columns. Difficult scenarios are involved, including occlusions, interfering targets, and small-scale pedestrians. Blue, red and purple bounding boxes stand for true positives, false positives, and false negatives, respectively. Corresponding ground truth is shown by green box for each true positive predictions.
4.3 Quantitative Metrics
The metric $\mathrm{MR}^{-2}$ is used for evaluating the accuracy on KAIST. It equals to the area of log-average Miss Rate (MR) over False Positive Per Image (FPPI), where FPPI ranges from $10^{-2}$ to $1$. Besides, the log-average MR-FPPI curves are obtained by varying the threshold of detector from 0 to $1$ to generate different MR-FPPI points and plotting these points in logarithmic-scale axes. For consistent comparison with previous works on FLIR, the COCO-style [COCO2014] evaluation metrics are adopted, namely Average Precision (AP) for specific class and mean Average Precision (mAP) for all classes. To evaluate the efficiency of multispectral pedestrian detectors, the memory cost are denoted by the numbers of parameters (Param.), while computational complexity are indicated by the number of floating point operations (FLOPs) and the inference time (Time). Notably, the inference time of our model includes post-processing time. Many previous works may not prioritize the efficiency of their models and have not report corresponding metrics in their papers, thus we estimate their lower bound. The FLOPs metric is estimated based on $640\times 640$ input size.